OpenStudy (anonymous):

3x(1)+(2). what is the ordered pair

OpenStudy (anonymous):

An ordered pair is a pair of numbers in a specific order.

OpenStudy (anonymous):

yeah, some guy earlier said to find the ordered pair and get back to me. so i did and hes not replying.

OpenStudy (anonymous):

oh :)

OpenStudy (anonymous):

haha, yeah

OpenStudy (anonymous):

is there a way to get the ordered pair from this?

OpenStudy (anonymous):

Ordered pairs are often used to represent two variables. for example When we write (x, y) = (7, - 2) , we mean x = 7 and y = - 2 . The number which corresponds to the value of x is called the x-coordinate and the number which corresponds to the value of y is called the y-coordinate.

OpenStudy (anonymous):

yeah, i know that part. im getting into substitution and i dont know a lick of of it. heres the origional problem......Solve this system of equations. Write your answer as an ordered pair. x + 3y = 1 -3x - 3y = -15

OpenStudy (anonymous):

in ur question there no ordered pair. hw can i solve ur question.

OpenStudy (anonymous):

could it be no solution

OpenStudy (anonymous):

without order pair no.

OpenStudy (anonymous):

iight, thank you....i have a few more problems after this one if you have time

OpenStudy (anonymous):

okay i ty :)

OpenStudy (anonymous):

Solve this system of equations. Write the answer as an ordered pair. -5x - 2y = -2 x - 7y = -7

OpenStudy (anonymous):

thats the next one